Bob Agrajag

Ability Card  Non-Transferable

Mop Combat

I hit you with my mop. Fall unconscious.

Hitchhiker’s Guide: B Ark  IN: Spring 1998

Bob Agrajag

Ability Card  Non-Transferable

Ability Effect:

You are a master of combat with a mop. This requires years of training, learning the balance of a particular mop. If you have your mop and are free to use it, you may knock out anyone in your ZOC instantly.

Hitchhiker’s Guide: B Ark  IN: Spring 1998